40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 19 05:54:50 2010, rana, Configuration, PSL, RC Temperature Servo Turned OFF temporarily Untitled.png
    Reply  Mon May 3 17:49:41 2010, rana, Configuration, PSL, RC Temperature Servo Turned OFF temporarily rct.png
       Reply  Tue May 4 02:28:38 2010, rana, Configuration, PSL, RC Temperature Servo Turned OFF temporarily 
          Reply  Tue May 4 06:32:58 2010, alberto, Configuration, PSL, RC Temperature Servo Turned OFF temporarily 
             Reply  Wed May 5 02:37:55 2010, rana, Configuration, PSL, RC Temperature Servo Turned OFF temporarily Picture_7.png
Message ID: 2881     Entry time: Wed May 5 02:37:55 2010     In reply to: 2876
Author: rana 
Type: Configuration 
Category: PSL 
Subject: RC Temperature Servo Turned OFF temporarily 

Quote:

 more detailed instructions needed....

I showed Kiwamu and Alberto how to turn the lights on and off in the PSL. This is why Caltech is such a fine institution: most schools would have TAs delivering this kind of optics instruction.

We've turned off the RC temperature stabilization and the lights will supply the quasi-random heat input to the table and the cavity. Alberto and Kiwamu will be turning the lights on and off at random times.

The attached plot is the spectrum of temperature fluctuations of the room and the vacuum can with no stabilization from this weekend. I think the rolloff above 10 mHz is kind of fake - I had the .SMOO parameter set to 0.99 for both of these channels. I've just now set the .SMOO to 0 for both channels, so we should now see the true ADC or sensor noise level. It should be ~1 mK/rHz.

Attachment 1: Picture_7.png  64 kB  | Hide | Hide all
Picture_7.png
ELOG V3.1.3-